OpenStudy (anonymous):

integral 1/-32-.015v dv= integral dt

OpenStudy (anonymous):

1/(-32-.015v) dv

OpenStudy (turingtest):

\[\int\frac1{-32-0.015v}dv\]this?

OpenStudy (turingtest):

\[u=-32-0.015v\implies du=?\]

OpenStudy (anonymous):

-.015 dv

OpenStudy (turingtest):

good, so\[du=-0.015dv\implies dv=?\]

OpenStudy (anonymous):

du/-.015

OpenStudy (turingtest):

great now plug that into the original intagral

OpenStudy (anonymous):

\[\int\limits_{?}^{?} 1/u du=\]

OpenStudy (anonymous):

-.015ln(-32-.015v)

OpenStudy (turingtest):

you are missing a constant....

OpenStudy (turingtest):

not quite, but close

OpenStudy (anonymous):

help please

OpenStudy (turingtest):

what was du again? you said it earlier...

OpenStudy (turingtest):

you are very close, but I want you to see your mistake

OpenStudy (anonymous):

-.015dv

OpenStudy (turingtest):

sorry I mean what was dv ? you said it earlier...

OpenStudy (turingtest):

yes, so dv=....?

OpenStudy (anonymous):

-1/.015 du

OpenStudy (turingtest):

good, so where did that -1/0.015 go when you made your substitution?

OpenStudy (turingtest):

you dropped it ;) gotta be a bit more careful there

OpenStudy (turingtest):

brb, try again now with that tip

OpenStudy (anonymous):

\[\int\limits_{?}^{?} 1/u (-1/.015)du=-1/.015 \ln u\]

OpenStudy (turingtest):

..+C yes, now just sub back in for u :)

OpenStudy (anonymous):

heres the problem: A woman bails out of an airplane at altitude of 10,000 ft falls freely for 20s, then opens parachute. How long will it take her to reach the ground? ssume linear air resistance pv ft/s^2, taking p=.15 without the parachute and p=1.5 with the parachute. (first determine her height aabove ground and velocity when the parachute opens) accelercation due to gravity would be -32 ft/s^2 so I dv/dt--32-.15v v<0 \[\int\limits_{?}^{?}1/(-32-.15v) dv=\int\limits_{?}^{?} dt\] give me -1/.15 ln(-32-.15t)= t+C v(0)= -1/.15 ln(-32-.15(0))=0+C c=?? since it says linear can I v=e^integral p = e^-.15t(integral -32e^-.15t)= e^-.15t(-32/.15 e^.15t+C) look at a solution and I do not follow so can you walk me through this?

OpenStudy (turingtest):

well now this is a word problem, so let me have lunch and I will get back to you on this, okay?

OpenStudy (anonymous):

thank you